OpenStudy (anonymous):

What is z in z = 163.35 * 45%?

OpenStudy (agreene):

z=163.35*0.45 z=73.5075

OpenStudy (anonymous):

163.35*0.45=73.5075

OpenStudy (anonymous):

Know the easiest method of reaching that answer?

OpenStudy (anonymous):

z=73.5075

OpenStudy (anonymous):

w/o a calc that is

OpenStudy (agreene):

using long multiplication would probably be the easiest way without a calculator.

OpenStudy (anonymous):

I see thanks